I need to know how to solve this problem. The question is Find the equation of the line that satisfies (2,3) Slope: 1
Answer by sarah_adam(201)

The general equation for a point , slope form is:
y = y1 + b (x - x1)
here (x1,y1) are the given coordintes and b is the slope
so here x1 = 2 and y1 = 3 and b = 1
substituting the values
y = 3 + 1(x-2)
y = 3 + x -2
y = 1 + 4x